A store has clearance items that have been marked down by 50%. They are having a sale, advertising an additional 40% off clearance items. What percent of the original price do you end up paying?
Answer by math_tutor2020(3835)


x = original price before any markdowns or sales or discounts
0.5x = price after the 50% markdown

When you have a 40% discount, then you pay the remaining 60%
60% of 0.5x = 0.6*0.5x = 0.3x
Effectively the two discounts combine to the customer paying 30% of the original price (meaning they get a 70% discount overall).

-----------------------------------------

An example:

Let's say an item is originally priced at $100
After the 50% markdown, it's now worth half as much at $50

If there's an additional 40% discount, then the customer saves 50*0.4 = 20 dollars and ends up paying 50-20 = 30 dollars when all is said and done (this doesn't include sales tax).

Note how the ratio of the final price (30) and the original (100) is 30/100 = 0.30 = 30%
Meaning the customer paid 30% of the starting price.

-----------------------------------------

Answer: 30%